Show that in a group G of odd order, the equation x2= e has a unique solution.

Further, show that x2= g has a unique solution ∀g∈G,g ≠e .


1
Expert's answer
2021-06-14T15:21:23-0400

The equation "x^2=g" will not have a unique solution if "\exist a,b \in G" such that "a^2=b^2" or "\nexists a \in G" such that "a^2=g." So "x^2=a" has a unique solution if and only if the function "f:G\to G" such that "f(x)=x^2" is bijective. Hence by (a), our result follows.